Solve the following pairs of simultaneous equations.
4x - y = 12
2y - 8x = -24

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

4 x -1 y = 12 .............1
Total value
-8 x + 2 y = -24 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
8 x -2 y = 24
-8 x + 2 y = -24
Add the two equations
0 x = 0
this is absurd
Both the equations are of the same line.
Hence there are infinite solutions

Answer by Alan3354(69443)

4x - y = 12
2y - 8x = -24 --> 4x - y = 12
------
It's the same eqn --> dependent, no unique solution.